# Machine Learners Should Acknowledge the Legal Implications of Large Language Models as Personal Data

### v1

@@ 3.1 Why LLMs Memorize Personal Data
Figure 1: LLMs and partially successful personal data editing techniques. (a) We aim to edit facts about personal data, in our case the president of the US, in GPT-J 6B with MEMIT (Meng et al., 2022, Mass-Editing Memory in a Transformer). (b) Pre-interventional model. The heatmap shows the influence of each word on the prediction across all layers. The test prompt yields the correct answer. We also include the response of an unrelated and related prompt to our intervention. Parentheses include the likelihood of the response. (c) Edited model. The test prompts yield the desired answer, the unrelated prompt remains unchanged, and the intervention-related prompt, however, still gives the unedited answer. The latter highlights the challenges of editing facts and data removal in LLMs. Icons from flaticon.com
@@
LLMs store training data in parameters through an interconnected and overlapping architecture. This complex structure makes it difficult to directly access or prove data memorization. But does personal data have to be human-perceptible? Under the GDPR, the format in which information is encoded is irrelevant when determining whether it qualifies as personal data (EDPB, 2024). For instance, JPEG, MP3, or PDF files are personal data if they allow an identification of an individual, in spite of their contents not being directly perceptible. Similarly, the fact that data is imperceptibly stored within AI model parameters does not preclude it from being considered personal data.  Furthermore, pseudonymization does not alter the status of personal data. Personal data that cannot be attributed to a specific individual without additional information is still regarded as personal data (Recital 26 GDPR). Personal data stored in AI models can be considered a form of pseudonymization (Veale et al., 2018).  Similarly, the imperceptibility of data stored in LLMs parameters does not exclude the possibility for personal data to be stored (EDPB, 2024). Moreover, just because we cannot directly access, observe, or pinpoint specific data within the model, this does not mean that it is not stored. To date, evidence of stored data in LLMs can only be inferred by observing correlations between inputs and outputs (Cooper & Grimmelmann, 2024). Studies estimate that LLMs can memorize 0.1 to 10 percent of their training data verbatim (Cooper & Grimmelmann, 2024). In addition, evidence suggests that larger models memorize more than smaller ones, and that data that is frequently repeated in the training set is more likely to be memorized (Lee et al., 2022; Carlini et al., 2023; Nasr et al., 2023b).
@@
Figure 1 shows a real-world example of GPT-J 6B storing personal data and model attribution. We used the tools from Meng et al. (2022) to edit the model. GPT-J 6B predicts the next word with the input “The last name of the president of the USA is” with the correct sequence “Trump”. Please note that we choose personal data that could also be considered common knowledge on purpose. We do not want to expose the personal data of a less-known individual. Our argument works with any personal data, whether the person is of public interest or not. The heatmap in Figure 1b also shows the attribution of different layers and the input sequence to the output prediction. Editing personal data will be discussed in Section 5.
@@
Some algorithms, like k-nearest neighbor classification and support vector machines, explicitly encode data points and make (personal) training data an integral part of the model (Brown et al., 2021). However, even neural networks have to balance memorization and generalization. In fact, research suggests some degree of memorization is essential for generalization (van den Burg & Williams, 2021; Chatterjee, 2018; Feldman & Zhang, 2020). Furthermore, compression through training reduces data but preserves essential details, much like a ZIP file (Cooper & Grimmelmann, 2024). As a result, even highly compressed models are likely to retain identifiable patterns or characteristics.
@@ 3.2 When Do LLMs Qualify as Personal Data?
Determining whether LLMs qualify as personal data under the GDPR depends on whether it is “reasonably likely” that an individual can be identified through the model’s output. This requires a case-by-case assessment and depends on the effort needed to retrieve personal data in the model (EDPB, 2024). A LLM can only be considered non-personal data, if no personal data from the training data can be extracted using reasonable means; and its outputs do not relate to the natural persons whose data was used to train it. This requires an impossibility to single out, link, and infer information from the supposedly anonymous LLM (EDPB, 2024)—a very high bar to clear.
@@ 3.3 Alternative Views
Our position is challenged by scholars and data protection authorities who argue that AI models, in general, cannot be classified as personal data (Wachter & Mittelstadt, 2019; Leiser & Dechesne, 2020; Datatilsynet, 2023; Hamburg, 2024). Consequently, they contend that the legal implications outlined in Section 4 do not apply. Their argument is that even specific AI models, such as LLMs, do not memorize data. Therefore, the argument goes, LLMs cannot, on their own, be considered personal data under the GDPR. Specifically, LLMs are compared to statistical reports, which are not regarded as personal data if they contain only conclusions and aggregated data derived from statistical analysis (Datatilsynet, 2023). It is further argued that personal data used to train LLMs is transformed into abstract mathematical representations and probability weights, which supposedly is not personal data (Hamburg, 2024). Instead of memorizing personal data, LLMs are said to merely learn correlations between tokens based on probability weights (Hamburg, 2024).
However, these arguments are flawed for the following four reasons: First, the format of the encoded information is irrelevant under the GDPR (see Section 1). Second, unlike a statistical report, an LLM can be queried, for instance, with an individual’s name to produce personal data. Third, while it is undisputed that the outputs of LLMs are probabilistic [^11] Even setting the “temperature” parameter of an LLM to zero does not lead to complete determinism (Ouyang et al., 2023). and often inaccurate (see Figure 2), data does not need to be fully accurate to be personal data under the GDPR. Even blurred images or partially incorrect details may suffice for identification, as fragments or partial replicas can reveal an individual’s identity depending on the context. Fourth, our experiments (see Figure 1) demonstrate that personal data is stored within LLMs and can even be edited. As a result, if personal data stored within an LLM can be extracted in a way that makes the identification of an individual “reasonably likely”, the LLM must be considered personal data under the GDPR (EDPB, 2024; Veale et al., 2018; Juliussen et al., 2023).
@@ 4 Overlooked Legal Implications of LLMs as Personal Data by the ML Community
Classifying an LLM as personal data carries profound legal implications for the machine learning community, potentially impacting data handling practices, compliance requirements, and model development processes. In the following, we will discuss different non-exhaustive legal requirements: First, researchers need a legal basis if they are training a model or providing the model on, e.g., GitHub or Hugging Face, such as consent, contract, or legitimate interest. Second, data subjects would have a right to access, delete, and rectify their personal data within the LLM. In the following, we focus on the regulations regarding a lawful basis, the right of access and deletion, and the obligation of data controllers to ensure privacy by design since these are, from our perspective, the most overlooked requirements in ML. Please note that we simplify some of the legal nuances for for clarity and due to space limitations, see (Feiler et al., 2018) for legal details.
@@
Article 89 GDPR states that the processing of personal data for scientific purposes shall be subject to appropriate safeguards. One of these safeguards is, for example, the principle of data minimization. From our point of view, this is hardly seen in LLM research; in contrast, more and more data are used (Bender et al., 2021; Shi et al., 2023; Kaplan et al., 2020). Article 89 of the GDPR also allows EU member states to provide derogations in the context of scientific research with regard to the rights of data subjects. However, the scope of Art. 89 is highly controversial (See (Kindt et al., 2021; Biega & Finck, 2021) for more details). Because of this controversial debate, we focus on the EU-wide law.
4.1 Lawful Basis: Researchers Require a Reason to Process Data
@@
Whenever an LLM is classified as personal data, any “processing” of the model requires a legal basis (see Section 2.3 for an explanation of the term processing; A full list of legal bases can be found in Article 6 GDPR). Therefore, whenever a model is shared, uploaded, downloaded, or distributed locally or on platforms such as GitHub or Hugging Face, these processes require a legal basis. Under the GDPR, the main legal bases for data processing include consent, the performance of a contract, protection of vital interests, and legitimate interests (Article 6(1) GDPR). These legal bases are complex and raise a number of legal issues, which we will not address here. Instead, we refer the interested reader to (Feiler et al., 2018).
@@
4.2 Right of Access: Researchers Have to Provide Individuals With Information About Their Personal Data Stored in LLMs
@@
According to Article 15 GDPR, data subjects have a right to obtain access to their personal data from the controller. This raises the question of how this right can be implemented, as it is almost impossible to immediately access specific data within the LLM, as mentioned above. It is unclear how data controllers can comply with such a request. They could either use structured prompts to query the model for potentially relevant personal data (Carlini et al., 2021), or ask data subjects to suggest their own prompts. Testing prompts under different conditions, similar to red teaming (Perez et al., 2022), and parameters can help identify data that is consistently included in the outputs, indicating a higher likelihood of personal data being stored. In addition, the information provided to the data subject would need to include a clarification that a complete extraction of all personal data from the model is not technically feasible, that the model may hallucinate, and that any answers should be presented as approximations.
@@
4.3 Right to Be Forgotten: At the Request of Data Subjects, Personal Data Must Be Deleted From LLMs
@@
Furthermore, data subjects have, under some circumstances, the right to request the erasure of their personal data from the controller within an LLM (Article 17 GDPR). However, fulfilling this obligation poses technical challenges (Veale et al., 2018; Villaronga et al., 2018; Cooper et al., 2024). Simply retraining the model is often not possible, keeping in mind that the pure energy costs for training an LLM can be several million dollars (de Vries, 2023). Different approaches to solve these issues have been discussed in ML Literature (Veale et al., 2018; Villaronga et al., 2018). Approaches for fast and easy “machine unlearning” (Bourtoule et al., 2021) have only recently been proposed and are still largely unexplored, let alone ready for use(Nguyen et al., 2022; Zhang et al., 2023b). The methods currently under discussion cannot be retrofitted into existing systems but would require a complete redesign of the entire model pipeline with unclear implications.
@@
4.4 Data Protection by Design: Researchers Are Required to Think About Data Protection From the Beginning
@@
In all cases, LLM developers must implement technical and organizational measures to comply with data protection principles (Article 25(1) GDPR). For example, developers must make their models resistant to LLM-specific privacy attacks (Yao et al., 2024; Li et al., 2023), such as data extraction attacks or membership inference attacks. Furthermore, Article 35(1) GDPR requires a data protection impact assessment for specific risky areas. Machine Learning, and specifically LLMs qualify as such a risky area (European Data Protection Supervisor, 2019). Developing this data protection impact assessment is a non-trivial task (Kloza et al., 2019). It usually requires critical thinking, extensive documentation and providing safeguards in advance.
4.5 Possible Legal Consequences for LLM Researchers
@@
The legal consequences of classifying an LLM as personal data under the GDPR are far-reaching. For example, if data controllers transfer a model without a legal basis to do so, and such a model is inverted, this would likely be considered both a data breach and a violation of the principle of security in general (Veale et al., 2018). In addition, breaches of obligations and rights of data subjects also trigger significant consequences. For particularly serious breaches, as defined in Art. 83(5) GDPR, the fine can be up to 20 million euros or, in the case of a company, up to 4 percent of its total worldwide turnover in the previous financial year, whichever is higher. But even the catalog of less serious breaches in Art. 83(4) GDPR provides for fines of up to 10 million euros or, in the case of a company, up to 2 percent of its total worldwide turnover in the previous financial year, whichever is higher. The fines are not merely a theoretical risk, as there have already been cases where individual researchers have faced penalties (APD, 2022) as well as at least 35 fines against universities (GDPR ET, 2024), see also (Ruohonen & Hjerppe, 2022) for more details.
4.6 Summary of Legal Implications of LLMs as Personal Data
@@ 5 Recommendation for Research Practice in the LLM Community
In this section, we provide brief recommendations for ML researchers in the LLM domain.
@@
The GDPR applies to the processing of personal data. If a model is trained without any personal data, it falls outside the scope of the GDPR. [^12] Note that in some rare cases, an LLM’s input data can generate personal data as output, even if that data was not stored in the model but was generated due to the input. For example, if a user asks the model to spellcheck a CV, the model will output personal data. The legal implications of these exceptional cases are beyond the scope of this paper. In this case, the GDPR does not apply. Researchers who train LLMs with personal data need to ensure that they have a legal Data used for training may require a legal basis or fall under a specific scientific exception. Even in the latter case, other obligations may still apply, such as the obligation for data minimization (Article 89(1) GDPR). It might also be possible to use Differential Privacy methods (Dwork, 2006) to avoid access of personal data by users of the model (Cummings & Desai, 2018; Holzel, 2019), see (Cummings et al., 2023) for more details.
@@
To prevent the exposure of personal data stored in an LLM, the model could be encapsulated in a privacy-preserving framework. Additional layers can be implemented to minimize the risk of disclosing stored personal data (Cooper & Grimmelmann, 2024). For example, at the “front end”, user input can be filtered or modified before being processed by the model. On the back end, models can be tuned to refuse to generate content that could lead to the re-identification of individuals. Finally, again on the “front end”, LLM outputs can be filtered or modified before being delivered to users. This layered architecture allows developers to mitigate the risk of exposing stored personal data. Consequently, it could be that some providers of LLMs may not need to remove specific data points from their trained model, but filter results at the output level.
@@
On a technical side, it might be possible to use and extend unlearning methods (Bourtoule et al., 2021; Nguyen et al., 2022) to comply with the right to be forgotten. The idea of machine unlearning is that it is possible to let the model “forget” specific data points. It might also be possible to use editing methods like MEMIT (Meng et al., 2022), which we used in the context of GPT-J 6B in Figure 1 to edit facts of personal data and thus remove them from the data. In Figure 1b, we edited the last name of the president of the USA from Trump to the president of France, “Macron”. After editing, the model outputs the correct information for the initial prompt. However, while unlearning tools in the context of LLMs pose a promising direction, there are still severe challenges (Xu et al., 2024). This also shows our example in Figure 1b. When using the “last name of the president of the USA”-unrelated prompt about the Eiffel Tower, the model answers unchanged. In contrast, asking it in a different way about the president of the US still results in the model memorizing the correct answer of “Trump”.
@@
Lastly, one might wonder whether the recently adopted EU Artificial Intelligence Act (AIA) [^13] EU Regulation 2024/1689, 12.7.2025. provides an answer on how to address the memorization of personal data in LLMs. The AIA aims to provide a legal framework for the development, deployment, and use of human-centered and trustworthy artificial intelligence (AI) [^14] Article 1(1) AIA.. It establishes legal requirements for specific types of AI systems and general-purpose AI models that must be fulfilled before they are placed on the market, inter alia, distributed or used in the EU in the course of a (commercial) activity. [^15] Article 3(10) AIA. Without going into detail, it is important to note that the AI Act applies without prejudice to the GDPR. [^16] Article 2(7) AIA. Therefore, both legal frameworks could apply simultaneously. However, it remains questionable whether the AI Act would be relevant to our topic and to ML researchers since the AI Act does not apply to AI systems or AI models, including their output, that are developed and put into service solely for the purpose of scientific research and development. [^17] Article 2(6) AIA. It is indeed questionable to what extent this legal exemption applies to ML researchers who publish their LLMs on public platforms. Therefore, the issue of memorization of personal data remains relevant only under the GDPR.
From our point of view, a combination of technical and compliance measures will be the favorable approach to protect personal data. Additionally, awareness in the ML community about challenges and solutions is beneficial.

### v2

@@ 3.1 Why LLMs Memorize Personal Data
Figure 1: LLMs and partially successful personal data editing techniques. (a) We aim to edit facts about personal data, in our case the president of the US, in GPT-J 6B with MEMIT [54, Mass-Editing Memory in a Transformer]. (b) Pre-interventional model. The heatmap shows the influence of each word on the prediction across all layers. The test prompt yields the correct answer. We also include the response to an unrelated and related prompt to our intervention. Parentheses include the likelihood of the response. (c) Edited model. The test prompts yield the desired answer, the unrelated prompt remains unchanged, and the intervention-related prompt, however, still gives the unedited answer. The latter highlights the challenges of editing facts and data removal in LLMs. Icons from flaticon.com
@@
LLMs store training data in parameters through an interconnected and overlapping architecture. This complex structure makes it difficult to directly access or prove data memorization. But does personal data have to be human-perceptible? Under the GDPR, the format in which information is encoded is irrelevant when determining whether it qualifies as personal data [24]. For instance, JPEG, MP3, or PDF files are personal data if they allow an identification of an individual, in spite of their contents not being directly perceptible. Similarly, the fact that data is imperceptibly stored within AI model parameters does not preclude it from being considered personal data.  Furthermore, pseudonymization does not alter the status of personal data. Personal data that cannot be attributed to a specific individual without additional information is still regarded as personal data (Recital 26 GDPR). Personal data stored in AI models can be considered a form of pseudonymization [76].  Similarly, the imperceptibility of data stored in LLMs parameters does not exclude the possibility for personal data to be stored [24]. Moreover, just because we cannot directly access, observe, or pinpoint specific data within the model, this does not mean that it is not stored. To date, evidence of stored data in LLMs can only be inferred by observing correlations between inputs and outputs [14]. Studies estimate that LLMs can memorize 0.1 to 10 percent of their training data verbatim [14]. In addition, evidence suggests that larger models memorize more than smaller ones, and that data that is frequently repeated in the training set is more likely to be memorized [50, 12, 56].
@@
Figure 1 shows a real-world example of GPT-J 6B storing personal data and model attribution. We used the tools from Meng et al. [54] to edit the model. GPT-J 6B predicts the next word with the input “The last name of the president of the USA is” with the correct sequence “Trump”. Please note that we choose personal data that could also be considered common knowledge on purpose. We do not want to expose the personal data of a less-known individual. Our argument works with any personal data, whether the person is of public interest or not. The heatmap in Figure 1b also shows the attribution of different layers and the input sequence to the output prediction. Editing personal data will be discussed in Section 5.
@@
Some algorithms, like k-nearest neighbor classification and support vector machines, explicitly encode data points and make (personal) training data an integral part of the model [7]. However, even neural networks have to balance memorization and generalization. In fact, research suggests some degree of memorization is essential for generalization [75, 13, 29]. Furthermore, compression through training reduces data but preserves essential details, much like a ZIP file [14]. As a result, even highly compressed models are likely to retain identifiable patterns or characteristics.
@@ 3.2 When Do LLMs Qualify as Personal Data?
Determining whether LLMs qualify as personal data under the GDPR depends on whether it is “reasonably likely” that an individual can be identified through the model’s output. This requires a case-by-case assessment and depends on the effort needed to retrieve personal data in the model [24]. A LLM can only be considered non-personal data, if no personal data from the training data can be extracted using reasonable means; and its outputs do not relate to the natural persons whose data was used to train it. This requires an impossibility to single out, link, and infer information from the supposedly anonymous LLM [24]—a very high bar to clear. This assessment also reflects the broader understanding that under the GDPR, the concept of “data” is to be interpreted as “information” about identifiable individuals, regardless of how it is stored or represented. In that sense, an LLM is not merely a processing system operating on a distinct dataset but rather constitutes a data source in its own right—one that embeds representations of personal data learned during training.
@@ 3.3 Alternative Views
Our position is challenged by scholars and data protection authorities who argue that AI models, in general, cannot be classified as personal data [81, 51, 20, 35]. Consequently, they contend that the legal implications outlined in Section 4 do not apply. Their argument is that even specific AI models, such as LLMs, do not memorize data. Therefore, the argument goes, LLMs cannot, on their own, be considered personal data under the GDPR. Specifically, LLMs are compared to statistical reports, which are not regarded as personal data if they contain only conclusions and aggregated data derived from statistical analysis [20]. It is further argued that personal data used to train LLMs is transformed into abstract mathematical representations and probability weights, which supposedly is not personal data [35]. Instead of memorizing personal data, LLMs are said to merely learn correlations between tokens based on probability weights [35].
However, these arguments are flawed for the following four reasons: First, the format of the encoded information is irrelevant under the GDPR (see Section 1). Second, unlike a statistical report, an LLM can be queried, for instance, with an individual’s name to produce personal data. Third, while it is undisputed that the outputs of LLMs are probabilistic [^12] Even setting the “temperature” parameter of an LLM to zero does not lead to complete determinism [59]. and often inaccurate (see Figure 2), even if parts of the personal data are not accurate, it is still considered personal data under the GDPR. This breaks down if the model hallucinates without any ground truth. Even blurred images or partially incorrect details may suffice for identification, as fragments or partial replicas can reveal an individual’s identity depending on the context. Fourth, our experiments (see Figure 1) demonstrate that personal data is stored within LLMs and can even be edited. As a result, if personal data stored within an LLM can be extracted in a way that makes the identification of an individual “reasonably likely”, the LLM must be considered personal data under the GDPR [24, 76, 45].
Recently, some papers discussed to what extent data memorization occurs in LLMs[41, 22, 14]. One could ask, whether LLMs memorize specific personal data. We believe that there is an important difference in perspective. One viewpoint has an individual focus: Are my personal data memorized in an LLM? In contrast, our perspective is that of the developer: Are there any personal data stored in the LLM? We note that for the applicability of the GDPR, it is sufficient that only a small fraction, indeed even one personal data item, is memorized. There is no minimal threshold for GDPR applicability. Thus, we think our paper is important to highlight the issues of LLMs as personal data for the ML community.
@@ 4 Overlooked Legal Implications of LLMs as Personal Data by the ML Community
Classifying an LLM as personal data carries profound legal implications for the machine learning community, potentially impacting data handling practices, compliance requirements, and model development processes. In the following, we will discuss different non-exhaustive legal requirements: First, researchers need a legal basis if they are training a model or providing the model on, e.g., GitHub or Hugging Face, such as consent, contract, or legitimate interest. Second, data subjects would have a right to access, delete, and rectify their personal data within the LLM. In the following, we focus on the regulations regarding a lawful basis, the right of access and deletion, and the obligation of data controllers to ensure privacy by design since these are, from our perspective, the most overlooked requirements in ML. Please note that we simplify some of the legal nuances for clarity and due to space limitations, see [28] for legal details.
We want to clarify that just because LLMs are considered personal data (in our view), it is not forbidden to use them. Rather, it simply triggers the applicability of the GDPR, and therefore one needs a legal basis, as well as compliance with all regulations laid out in this section of our paper.
@@
Article 89 GDPR states that the processing of personal data for scientific purposes shall be subject to appropriate safeguards. One of these safeguards is, for example, the principle of data minimization. From our point of view, this is hardly seen in LLM research; in contrast, more and more data are used [3, 71, 46]. Article 89 of the GDPR also allows EU member states to provide derogations in the context of scientific research with regard to the rights of data subjects. However, the scope of Art. 89 is highly controversial (See [48, 5] for more details). Because of this controversial debate, we focus on the EU-wide law.
Lawful Basis: Researchers Require a Reason to Process Data
@@
Whenever an LLM is classified as personal data, any “processing” of the model requires a legal basis (see Section 2.3 for an explanation of the term processing; A full list of legal bases can be found in Article 6 GDPR). Therefore, whenever a model is shared, uploaded, downloaded, or distributed locally or on platforms such as GitHub or Hugging Face, these processes require a legal basis. Under the GDPR, the main legal bases for data processing include consent, the performance of a contract, protection of vital interests, and legitimate interests (Article 6(1) GDPR). These legal bases are complex and raise a number of legal issues, which we will not address here. Instead, we refer the interested reader to [28].
@@
Right of Access: Researchers Have to Provide Individuals With Information About Their Personal Data Stored in LLMs
@@
According to Article 15 GDPR, data subjects have a right to obtain access to their personal data from the controller. This raises the question of how this right can be implemented, as it is almost impossible to immediately access specific data within the LLM, as mentioned above. It is unclear how data controllers can comply with such a request. They could either use structured prompts to query the model for potentially relevant personal data [10], or ask data subjects to suggest their own prompts. Testing prompts under different conditions, similar to red teaming [61], and parameters can help identify data that is consistently included in the outputs, indicating a higher likelihood of personal data being stored. In addition, the information provided to the data subject would need to include a clarification that a complete extraction of all personal data from the model is not technically feasible, that the model may hallucinate, and that any answers should be presented as approximations.
@@
Right to Be Forgotten: At the Request of Data Subjects, Personal Data Must Be Deleted From LLMs
@@
Furthermore, data subjects have, under some circumstances, the right to request the erasure of their personal data from the controller within an LLM (Article 17 GDPR). However, fulfilling this obligation poses technical challenges [76, 79, 15]. Simply retraining the model is often not possible, keeping in mind that the pure energy costs for training an LLM can be several million dollars [21]. Different approaches to solve these issues have been discussed in ML Literature [76, 79]. Approaches for fast and easy “machine unlearning” [6] have only recently been proposed and are still largely unexplored, let alone ready for use[57, 90]. The methods currently under discussion cannot be retrofitted into existing systems but would require a complete redesign of the entire model pipeline with unclear implications.
@@
Data Protection by Design: Researchers Are Required to Think About Data Protection From the Beginning
@@
In all cases, LLM developers must implement technical and organizational measures to comply with data protection principles (Article 25(1) GDPR). For example, developers must make their models resistant to LLM-specific privacy attacks [87, 52], such as data extraction attacks or membership inference attacks. Furthermore, Article 35(1) GDPR requires a data protection impact assessment for specific risky areas. Machine Learning, and specifically LLMs qualify as such a risky area [26]. Developing this data protection impact assessment is a non-trivial task [49]. It usually requires critical thinking, extensive documentation and providing safeguards in advance.
Possible Legal Consequences for LLM Researchers
@@
The legal consequences of classifying an LLM as personal data under the GDPR are far-reaching. For example, if data controllers transfer a model without a legal basis to do so, and such a model is inverted, this would likely be considered both a data breach and a violation of the principle of security in general [76]. In addition, breaches of obligations and rights of data subjects also trigger significant consequences. For particularly serious breaches, as defined in Art. 83(5) GDPR, the fine can be up to 20 million euros or, in the case of a company, up to 4 percent of its total worldwide turnover in the previous financial year, whichever is higher. But even the catalog of less serious breaches in Art. 83(4) GDPR provides for fines of up to 10 million euros or, in the case of a company, up to 2 percent of its total worldwide turnover in the previous financial year, whichever is higher. The fines are not merely a theoretical risk, as there have already been cases where individual researchers have faced penalties [2] as well as at least 35 fines against universities [32], see also [68] for more details.
Summary of Legal Implications of LLMs as Personal Data
@@ 5 Recommendation for Research Practice in the LLM Community
The GDPR applies to the processing of personal data. If a model is trained without any personal data, it falls outside the scope of the GDPR [^13] Note that in some rare cases, an LLM’s input data can generate personal data as output, even if that data was not stored in the model but was generated due to the input. For example, if a user asks the model to spellcheck a CV, the model will output personal data. The legal implications of these exceptional cases are beyond the scope of this paper.. In this case, the GDPR does not apply. Researchers who train LLMs with personal data need to ensure that they have a legal basis. Thus, training may require a legal basis or fall under a specific scientific exception. Even in the latter case, other obligations may still apply, such as the obligation for data minimization (Article 89(1) GDPR). It might also be possible to use Differential Privacy methods [23] to avoid access to personal data by users of the model [18, 38], see [19] for more details.
@@
To prevent the exposure of personal data stored in an LLM, the model could be encapsulated in a privacy-preserving framework. Additional layers can be implemented to minimize the risk of disclosing stored personal data [14]. For example, at the “front end”, user input can be filtered or modified before being processed by the model. On the back end, models can be tuned to refuse to generate content that could lead to the re-identification of individuals. Finally, again on the “front end”, LLM outputs can be filtered or modified before being delivered to users. This layered architecture allows developers to mitigate the risk of exposing stored personal data. Consequently, it could be that some providers of LLMs may not need to remove specific data points from their trained model, but filter results at the output level.
@@
On a technical side, it might be possible to use and extend unlearning methods [6, 57] to comply with the right to be forgotten. The idea of machine unlearning is that it is possible to let the model “forget” specific data points. It might also be possible to use editing methods like MEMIT [54], which we used in the context of GPT-J 6B in Figure 1 to edit facts of personal data and thus remove them from the data. In Figure 1b, we edited the last name of the president of the USA from Trump to the president of France, “Macron”. After editing, the model outputs the correct information for the initial prompt. However, while unlearning tools in the context of LLMs pose a promising direction, there are still severe challenges [85]. This also shows our example in Figure 1b. When using the “last name of the president of the USA”-unrelated prompt about the Eiffel Tower, the model answers unchanged. In contrast, asking it in a different way about the president of the US still results in the model memorizing the correct answer of “Trump”.
@@
Lastly, one might wonder whether the recently adopted EU Artificial Intelligence Act (AIA) [^14] EU Regulation 2024/1689 provides an answer on how to address the memorization of personal data in LLMs. The AIA aims to provide a legal framework for the development, deployment, and use of human-centered and trustworthy artificial intelligence. It establishes legal requirements for specific types of AI systems and general-purpose AI models that must be fulfilled before they are placed on the market, inter alia, distributed or used in the EU in the course of a (commercial) activity [^15] Article 3(10) AIA.. Without going into detail, it is important to note that the AI Act applies without prejudice to the GDPR [^16] Article 2(7) AIA.. Therefore, both legal frameworks could apply simultaneously. However, it remains questionable whether the AI Act would be relevant to our topic and to ML researchers since the AI Act does not apply to AI systems or AI models, including their output, that are developed and put into service solely for the purpose of scientific research and development [^17] Article 2(6) AIA. It is indeed questionable to what extent this legal exemption applies to ML researchers who publish their LLMs on public platforms.. Therefore, the issue of memorization of personal data remains relevant only under the GDPR. From our point of view, a combination of technical and compliance measures will be the favorable approach to protect personal data. Additionally, awareness in the ML community about challenges and solutions is beneficial.